# The role of porosity in the transition to inertial regime in porous media flows

### v1

@@ 2.3 Post processing of the results
The initial positions of the tracers depend on the task at hand. In the case of $\phi=0.1$ SC sample, for the calculation of the streamline-based tortuosity $T_{s}$, we place them uniformly at the fluid part of the $x_{0}=0$ plane, with the distance $h_{s}=10^{-3}$ between the tracers along $y$- and $z$-direction. To find the recirculation/percolating volumes, we place the tracers only in the vicinity of the solid boundaries at $x_{0}=0$ plane, with the distance $h_{s}=10^{-3}$ between the tracers. In both cases, streamlines that do not connect the inlet and the outlet, hit the solid wall, or have negative increment in $x_{0}$-direction are not included in the further analysis.
@@
In case of $\phi=0.999$ SC sample, for the calculation of $T_{s}$ we place the tracers on a square grid of step $h_{s}=10^{-2}$ at the $x_{0}=0$ plane and dismiss the streamlines that do not fulfill the previously stated criteria. Due to numerical errors, the mean values of $x_{1}$- and $x_{2}$-velocities in this case are not identically zero, although they are two orders of magnitude lower than the mean $x_{0}$-velocity. It does not impact the calculation of $T_{\Omega}$ significantly, however causes the nearly-straight streamlines to experience a drift in the direction perpendicular to $x_{0}$. This results in their elongation form $L$ by a factor of magnitude comparable to the value of $T_{s}-1$. Due to this, for the higher-porosity SC sample, in Eq. (6) we use the distance between the streamline’s start and end as the normalization length, rather than the porous sample’s side length $L$. For the determination of the recirculation/percolating volume we use two sets of the initial positions of the tracers. To properly capture the flow separation on the rear part of the sphere’s surface we place the tracers upwind from the sphere, in the vicinity of the domain’s centerline $x_{1}=x_{2}=1$. After the integration, we apply the same criteria to dismiss certain streamlines as previously. To properly capture the saddle point at the end of the recirculation zone in the wake of the sphere, we place the tracers downwind from the sphere (also in the vicinity of the centerline) and integrate them in the upwind direction. We do not apply the previously described streamline exclusion criteria, rather we stop the integration before the streamlines reach the separation ring on the sphere’s surface, to avoid them falling into the recirculation zone due to the numerical errors. We use the separation $h_{s}=5\cdot 10^{-4}$ between the streamline starting positions along $y$- and $z$-direction for the identification of the recirculation/percolating volume.
Determining the recirculation and percolating volumes of fluid
@@
To determine which points of the fluid domain $\Omega$ of the SC samples belong to the recirculation or percolating volume we divide the periodic cell into $N_{H}^{3}$ cubic cells and consider only those occupied by the fluid. We then mark those of the cells which contain at least one point of the streamlines generated for this purpose. Having obtained the cluster of the marked cells, we consider those lying on the cluster’s boundary (i.e. having at least one non-marked neighbor), near the expected position of the boundary between the recirculation and percolating volume, as the approximation of this boundary. Lastly, all the cells lying on either side of the approximate boundary are marked as belonging to the recirculation or the percolating volume. We use the values of $N_{H}=160$ for both systems.
Calculation of $T_{\Omega}$ constituents
@@
To calculate the integrals of $V_{0}$ and $V$ over the whole volume of fluid, present in the definition of the volume-inegrated tortuosity $T_{\Omega}$, in the case of stochastic PM, the first-order method, i.e. summing the values of the hydrodynamic fields from each lattice point, is used. In the case of SC samples we use the ParaView utility IntegrateVariables [37, 38]. For the integration of $V_{0}$ and $V$ in the recirculation or the percolating volume we use the first order method, where the values of the fields are interpolated to centers of the indicator function using the same RBF-FD method as for the generation of streamlines.
@@ 3 Results
Difference in tortuosity–Reynolds number relation for porous samples of different porosities.
@@
We first investigate the transition from the Darcy to inertial regime for two stochastic porous media (PM) samples of porosities $\phi=0.7$ and $\phi=0.9$. Fig. 4 shows the dependence of the tortuosity $T_{\Omega}$ defined in Eq. (8) on the Reynolds number
$Re=\frac{\sqrt{k_{0}}\langle V_{0}\rangle}{\nu\phi},\quad\langle V_{0}\rangle=\frac{I[V_{0}]}{I[1]}$ (11)
@@
where $k_{0}$ is the permeability at $Re_{k}\rightarrow 0$. We note that for $Re_{k}$ values beyond the investigated range the solution went unsteady (transient). We observe that in the Stokes regime the tortuosity for each sample have a constant value, $1.07$ for $\phi=0.9$ and $1.175$ for $\phi=0.7$. As $Re$ increases, the tortuosity first decreases and then increases for both samples. However, the magnitude of the initial decrease and subsequent increase is considerably different for the two porosities. In the case of higher-porosity sample, those the decrease is about $0.02$ ($2/7$ of the initial value above unity) and the increase is about $0.01$. For the lower-porosity sample, these values are $0.003$ ($1/60$ of the initial value above unity) and $0.125$ ($0.7$ of the initial value above unity), respectively. Such differences between the minimal and maximal values achieved by $T_{\Omega}$ in the steady-state regime were previously observed in stochastic high-porosity media [12] and in low-porosity geological samples [14]. The onset of inertial effects is visible also in the visualizations of the velocity fields, as the weaker channeling at the outlets of the samples (Fig. 5) and the emergence of the recirculation zones in the wakes of the obstacles (Fig. 6).
@@
Figure 4: $T_{\Omega}$ for stochastic porous media of porosity $\phi=0.7$ (left) and $\phi=0.9$ (right). The horizontal dashed lines denote the values of $T_{\Omega}$ in the limit $Re_{k}\rightarrow 0$. The inset in the left subplot shows the same data with different ranges of the axes.
@@
Figure 5: Streamlines of the velocity field for the stochastic porous media of porosity $\phi=0.7$ (top row) and $\phi=0.9$ (bottom row) at low (left column) and high (right column) $Re_{k}$. The flow is along $x$ axis. The shown streamlines are generated using the streamTracer utility of ParaView.
@@
Figure 6: Velocity vectors at a chosen sections for the stochastic porous media of porosity $\phi=0.7$ (top row) and $\phi=0.9$ (bottom row) at low (left column) and high (right column) $Re_{k}$. The flow is from left to right. For $\phi=0.7$, only every 3rd velocity vector in each direction is shown for clarity.
Splitting the volume-integrated tortuosity $T_{\Omega}$
@@
In order to explain the difference between the tortuosity – Reynolds number relation described above, we first take a closer look at the definition of $T_{\Omega}$ and observe that both the numerator and denominator can be split into separate contributions coming from the percolating volume (denoted by subscript $p$) and the vortex, or recirculation, volume (denoted by subscript $v$)
@@
According to the definition of the vortex from Sec. 2, the mean velocity of the vortex in steady-state flow is zero, i.e. $I[\boldsymbol{V}]_{v}=\boldsymbol{0}$, so the term $I[V_{0}]_{v}$ in the denominator can be omitted. Thus, we can rewrite Eq. (12) as
@@
As the contribution of the recirculating zones to the value of the streamline-based tortuosity $T_{s}$ vanishes (cf. Eq. (7)), using the same procedure as in [25], it is possible to show that the integrals over the intersection of the surface $A$ and the fluid volume $\Omega$ in Eq. (7) can be transformed into the integrals over the percolating volume $\Omega_{p}$, i.e. $T_{s}=I[V]_{p}/I[V_{0}]_{p}$. This allows to rewrite Eq. (13) as
@@
Such form of the volume-integrated tortuosity allows to clearly distinguish between the contributions from the percolating and the recirculation zones. The denominator, $I[V_{0}]_{p}$, is proportional to the Reynolds number (cf. Eq. (11)) and does not introduce any new unknown variable into our considerations.
@@
As noted previously, the proper identification of the percolating and recirculating volumes in the fluid is very sensitive to stray streamlines (i.e. those that due to numerical errors cross the boundary between the percolating and recirculating volumes). During the present research we observed that even several stray streamlines can significantly alter the value of the streamline-based tortuosity. Because of this, we conduct the numerical investigation of the two contributions to $T_{\Omega}$ on a simple cubic (SC) systems and use thus obtained results to explain the behavior of $T_{\Omega}$ in the stochastic PM. SC samples can be thought of as models of particular fragments of a stochastic PM, at the same time having the geometry simple enough to allow for the accurate calculation of streamlines and identification of $\Omega_{p}$ and $\Omega_{v}$. To include in our analysis the cases with extremely high and low geometrical confinement, we consider SC systems of porosities $\phi=0.1$ (obstacle radius $0.6526$) and $\phi=0.999$ (obstacle radius $0.062$). Exemplary visualizations of the percolating volume in SC systems are shown in Fig. 8.
@@
Figure 7: Visualizations of the percolating volumes in $\phi=0.1$ SC system for low (left) and high (right) $Re_{k}$. The gray surfaces are the obstacles’ walls and for the low-porosity system only a few chosen walls are shown for clarity. The visualizations use the histograms from which the percolating volumes were found. We show only the percolating volumes for clarity.
Figure 8: Visualizations of the recirculating volumes for $\phi=0.999$ SC system for intermediate (left) and high (right) $Re_{k}$ (for low $Re_{k}$ the recirculation zone is not present). The gray surfaces are the obstacles’ walls and for the low-porosity system only a few chosen walls are shown for clarity. The visualizations use the histograms from which the recirculation volumes were found. We show only the recirculation volumes for clarity.
Tortuosity as the inertia onset indicator in simple cubic (SC) systems
@@
We begin our analysis by looking at the Reynolds number dependence of $T_{\Omega}$ and $T_{s}$ for SC sample, shown in Fig. 9. In the Darcy regime ($Re_{k}\rightarrow 0$), the values of both tortuosities stay constant and are compliant with each other, as shown in [25]. When the Reynolds number increases, the streamline-based tortuosity, $T_{s}$, decreases for both systems, which was observed in stochastic and periodic porous media by other authors [20, 13]. On the other hand, when the volume-based tortuosity, $T_{\Omega}$, is concerned, the inertial effects manifests themselves differently depending on the porosity, similarly to the stochastic porous media (cf. Fig. 4). In the lower-porosity case, its value decreases slightly in the range $Re_{k}\in[10^{-3};10^{-2}]$ to later increase significantly until the steady-state regime ends. Such behavior resembles the one observed for the lower-porosity stochastic porous media in Fig. 4 and also for a real-life porous sample [14]. For higher-porosity sample, the volume-based tortuosity follows the streamline-based one closely, to diverge from it only at $Re_{k}\approx 10^{-3}$. Similarly to the lower-porosity system, the initial decline of $T_{s}$ is followed by its rise, however to a much smaller extent. This is in line with the behavior of $T_{\Omega}$ for the higher-porosity stochastic porous media shown in Fig. 4, as well as a similar system investigated previously [12].
@@
The strictly decreasing behavior of the streamline-based tortuosity can be attributed to the rectification of streamlines which follow the growth of the recirculation zones volumes in the cavities and obstacles wakes. The volume-integrated tortuosity exhibits a more complex behavior, dependent on the sample’s porosity. Eq. (14) shows that the difference between $T_{\Omega}$ and $T_{s}$ is related to the integral of velocity magnitude inside the recirculation zones. Thus, we will further investigate the fall and the rise of the volume-based tortuosity in terms of the volume of the recirculation zones and the field of the velocity magnitude therein.
@@
Origin of the $T_{\Omega,v}$ decrease and rise – volume of the recirculation zones
Let us first focus on the numerator of the percolation-volume related term from Eq. (14), $I[V]_{v}/I[V_{0}]_{p}$. For convenience, we divide it by the total fluid volume, $I[1]$, i.e.

### v2

@@ 2.3.1 Generation of streamlines
The initial positions of the tracers depend on the task at hand. Due to the very different geometries and steady-state velocity fields in each of the SC samples, we tailor-cut the streamline generation for each case. For $\phi=0.1$ SC sample, for the calculation of the streamline-based tortuosity $T_{s}$, we place them uniformly at the fluid part of the $x_{0}=0$ plane, with the distance $h_{s}=10^{-3}$ between the tracers along $x_{1}$- and $x_{2}$-direction (see Fig. 8). To find the recirculation/percolating volumes, we place the tracers only in the vicinity of the solid boundaries at $x_{0}=0$ plane, with the distance $h_{s}=10^{-3}$ between the tracers. For this task, the grid from which the positions of the tracers are taken is rotated by $\pi/4$ about $x_{0}$-axis for a better compliance with the solid boundaries. For both tasks, we use the timestep length $\Delta t_{s}=10^{-3}$ during the integration along characteristics. Considering the initial positions of the streamlines for both tasks, theoretically, none of them should hit the solid wall, have a negative increment in $x$-direction, and all the streamlines should connect the inlet and the outlet (should be percolating). If any streamline fails to satisfy the mentioned conditions, we assume this is due to numerical errors and exclude it from further analysis.
Figure 8: Initial positions of the streamlines for the calculation of $T_{s}$ (left) and the identification of the recirculation/percolating volume (middle) for the low-porosity SC sample. The rightmost image shows which part of the domain is shown in the remaining ones. The fluid volume $\Omega$ is depicted in blue.
@@
In case of $\phi=0.999$ SC sample, for the calculation of $T_{s}$ we place the tracers on a square grid of step $h_{s}=10^{-2}$ at the $x_{0}=0$ plane (see Fig. 9, top-left) and use the timestep length $\Delta t_{s}=10^{-2}$. We dismiss the streamlines that do not fulfill the previously stated criteria. Due to numerical errors in the solution of the governing equations, the mean values of $x_{1}$- and $x_{2}$-velocities in this case are not identically zero, although they are two orders of magnitude lower than the mean $x_{0}$-velocity. It does not impact the calculation of $T_{\Omega}$ significantly, however, it causes the nearly-straight streamlines to experience a drift in a direction perpendicular to $x_{0}$. This results in their elongation form $L$ by a certain value, the magnitude of which is comparable to the value of $T_{s}-1$. In turn, the streamline-based tortuosity is artificially increased to a non-negligible extent. Due to this, for the higher-porosity SC sample, in Eq. (6) we use the distance between the streamline’s start and end as the normalization length, rather than the porous sample’s side length $L$. For the determination of the recirculation/percolating volume, we use two sets of the initial positions of the tracers (see Fig. 9, top-right). To properly capture the flow separation on the rear part of the sphere’s surface, we place the tracers upwind from the sphere, in the vicinity of the domain’s centerline $x_{1}=x_{2}=0.5$ (see Fig. 9, bottom-left). After the integration, we apply the same criteria to dismiss certain streamlines as previously. To properly capture the saddle point at the end of the recirculation zone in the wake of the sphere, we place the tracers downwind from the sphere (also in the vicinity of the centerline) and integrate them in the upwind direction (see Fig. 9, bottom-right). We do not apply the previously described streamline exclusion criteria; rather,, we stop the integration before the streamlines reach the separation ring on the sphere’s surface, to avoid them falling into the recirculation zone due to the numerical errors. We use the separation $h_{s}=5\cdot 10^{-4}$ between the streamline starting positions along $x_{1}$- and $x_{2}$-direction and the timestep length $\Delta t_{s}=10^{-3}$ for the identification of the recirculation/percolating volume.
Figure 9: Initial positions of the streamlines for the calculations in the high-porosity SC sample. Top-left: for the calculation of $T_{s}$. Top-right: for the identification of recirculation/percolating volumes. Bottom row: magnified views at the initial positions of the forward- (A) and backward-integrated (B) streamlines. The $x_{1}=x_{2}=0.5$ line is shown for clarity.
2.3.2 Determining the recirculation and percolating volumes of fluid
@@
To determine which points of the fluid domain $\Omega$ of the SC samples belong to the recirculation or percolating volume, the crucial part is to approximate the boundary between the two, $\partial\Omega_{pv}$. To do this, first we divide the domain into $N_{H}^{3}$ cubic cells and consider only those occupied by the fluid. We use $N_{H}=160$ for both systems. We then mark those of the cells which contain at least one point of the streamlines generated for the purpose of identification of $\Omega_{p}$ and $\Omega_{v}$ (Fig. 10, left). As mentioned above, we generate only the percolating streamlines (although not all of them). Due to this, holes may appear in the clusters of marked cells, denoted by an asterisk. Having obtained the cluster of the marked cells, we consider those lying on the cluster’s boundary (i.e. having at least one non-marked neighbor), near the expected position of $\partial\Omega_{pv}$, as the approximation of $\partial\Omega_{pv}$ (Fig. 10, middle). Lastly, all the cells lying on either side of the approximate boundary are marked as belonging to the recirculation or the percolating volume. The last step of this procedure automatically takes care of the proper coloring of the cells which do not contain any streamlines but should do so, if sufficiently many percolating streamlines were generated (the ones marked with an asterisk in Fig. 10). This way, one needs to generate the streamlines only in the vicinity of $\partial\Omega_{pv}$.
Figure 10: Schematic representation of the procedure for identifying the percolating ($\Omega_{p}$) and recirculation ($\Omega_{v}$) volumes. Each square represents a cell into which we divide the fluid volume. We note that those cells do not comply with the discretization. The percolating streamlines are plotted with thick dashed lines. We refer the reader to the paragraph Determining the recirculation and percolating volumes of fluid in Sec. 2.3 for the detailed explanation of the identification procedure.
2.3.3 Calculation of $T_{\Omega}$ constituents
@@
To calculate the integrals of $V_{0}$ and $V$ over the whole volume of fluid in Eq. (8), in the case of stochastic porous samples, the first-order method, i.e., the summing of the values of the hydrodynamic fields from each lattice point, is used. In the case of SC samples, we use the ParaView utility IntegrateVariables [42, 43]. For the integration of $V_{0}$ and $V$ in the recirculation or the percolating volume in SC samples, we interpolate $V$ and $V_{0}$ to the centers of the cells used for the identification of $\Omega_{p}$ and $\Omega_{v}$. We use the same RBF-FD interpolation as in the generation of streamlines. Finally, we use the first-order method for the integration of the fields in each of the volumes.
@@ 3 Results
3.1 Difference in tortuosity–Reynolds number relation for porous samples of different porosities.
@@
We first investigate the transition from the Darcy to inertial regime for two stochastic porous samples of porosities $\phi=0.7$ and $\phi=0.9$. Left subplot of Fig. 11 shows the dependence of the tortuosity $T_{\Omega}$ defined in Eq. (8) on the Reynolds number
$Re_{k}=\frac{\sqrt{k(Re_{k}\rightarrow 0)}\langle V_{0}\rangle\phi}{\nu},\quad\langle V_{0}\rangle=\frac{I[V_{0}]}{I[1]}$ (11)
@@
where $k(Re_{k}\rightarrow 0)$ is the permeability at the vanishing Reynolds number. The same data are shown rescaled in the right subplot of Fig. 11, where by $T_{\Omega}(Re_{k}\rightarrow 0)$ we mean the value of $T_{\Omega}$ at the vanishing Reynolds number (denoted by the horizontal dashed line in the left subplot). We note that for $Re_{k}$ values beyond the investigated range, the solution went unsteady. We observe that in the Stokes regime, the tortuosity for each sample has a constant value, $1.07$ for $\phi=0.9$ and $1.175$ for $\phi=0.7$. As $Re_{k}$ increases, the tortuosity first decreases and then increases for both samples. However, the magnitude of the initial decrease and subsequent increase is considerably different for the two porosities. In the case of the higher-porosity sample, the difference between the minimum of $T_{\Omega}$ and the $T_{\Omega}(Re_{k}\rightarrow 0)$ is about $0.02$ ($0.3$ relative to the Darcy-regime value above unity, see the right subplot), and the increase is about $0.01$. For the lower-porosity sample, these values are $0.003$ ($0.01$ relative to the Darcy-regime value above unity) and $0.125$ ($0.67$ relative to the Darcy-regime value above unity), respectively. Such differences between the minimal and maximal values achieved by $T_{\Omega}$ in the steady-state regime were previously observed in stochastic high-porosity media [14] and in low-porosity geological samples [16]. The onset of inertial effects is also visible in the visualizations of the velocity fields, as the weaker channeling at the outlets of the samples (Fig. 12) and the emergence of the recirculation zones in the wakes of the obstacles (Fig. 13).
@@
Figure 11: Left: the values of $T_{\Omega}$ for stochastic porous media of porosity $\phi=0.7$ (circles) and $\phi=0.9$ (squares) during the Darcy-inertial transition. The horizontal dashed lines denote the values of $T_{\Omega}$ in the limit of vanishingly small $Re_{k}$ ($Re_{k}\rightarrow 0$). Right: the same data as in the left subplot but rescaled by $T_{\Omega}$ at vanishingly small $Re_{k}$.
@@
Figure 12: Streamlines of the velocity field for the stochastic porous media of porosity $\phi=0.7$ (top row) and $\phi=0.9$ (bottom row) at low (left column) and high (right column) $Re_{k}$. The flow is along $x_{0}$ axis. The shown streamlines are generated using the streamTracer utility of ParaView for the sake of the visualization only and are not used for the calculation of $T_{s}$ or the identification of percolating/recirculation volumes.
@@
Figure 13: Velocity vectors at chosen sections for the stochastic porous media of porosity $\phi=0.7$ (top row) and $\phi=0.9$ (bottom row) at low (left column) and high (right column) $Re_{k}$. The flow is from left to right. For $\phi=0.7$, only every 3rd velocity vector in each direction is shown for clarity.
3.2 Splitting the volume-integrated tortuosity $T_{\Omega}$
@@
To explain such a different behavior of $T_{\Omega}(Re_{k})$ in the low- and high-porosity stochastic systems, we first take a closer look at the definition of $T_{\Omega}$ and observe that both the numerator and denominator can be split into separate contributions coming from the percolating volume $\Omega_{p}$ and the recirculation volume $\Omega_{v}$
@@
According to the definition of the vortex from Sec. 2, the mean velocity of the vortex in steady-state flow is zero, i.e., $I[\boldsymbol{V}]_{v}=\boldsymbol{0}$, so the term $I[V_{0}]_{v}$ in the denominator can be omitted. Thus, we can rewrite Eq. (12) as
@@
Now, using the same procedure as in [27], it is possible to show that the surface integrals over $A\cap\Omega_{p}$ in Eq. (7) can be transformed into volume integrals over the entire percolating volume $\Omega_{p}$, i.e., $T_{s}=I[V]_{p}/I[V_{0}]_{p}$. This allows to rewrite Eq. (13) as
@@
Such a form of the volume-integrated tortuosity allows one to clearly distinguish between the contributions from the percolating and the recirculation zones and directly points to the sources of inequality in Eq. (10). The denominator, $I[V_{0}]_{p}$, is proportional to the Reynolds number (cf. Eq. (11)) and does not introduce any new unknown variable into our considerations.
@@
As noted previously, the proper identification of the percolating and recirculating volumes in the fluid is very sensitive to stray streamlines (i.e., those that, due to numerical errors, cross the boundary between the percolating and recirculating volumes). We observed that even several stray streamlines can significantly alter the value of the streamline-based tortuosity. Because of this, we conduct the numerical investigation of the two contributions to $T_{\Omega}$ on simple cubic (SC) systems, and use the obtained results to explain the behavior of $T_{\Omega}$ in the stochastic samples. SC samples can be thought of as models of particular fragments of a stochastic PM, at the same time having the geometry simple enough to allow for the accurate calculation of streamlines and identification of $\Omega_{p}$ and $\Omega_{v}$. To include in our analysis the cases with extremely high and low geometrical confinement, we consider SC systems of porosities $\phi=0.1$ and $\phi=0.999$. Exemplary visualizations of the steady-state velocity fields in the studied SC systems are shown in Figs. 14 and 15.
@@
Figure 14: Visualizations of the sections of the velocity field in $\phi=0.1$ SC system for $Re_{k}=2.61\cdot 10^{-7}$ (left) and $Re_{k}=0.049$ (right). The gray surfaces are the obstacles’ walls, and only a few chosen walls are shown for clarity.
Figure 15: Visualizations of the sections of the velocity field in $\phi=0.999$ SC system for $Re_{k}=376$ (left) and $Re_{k}=1383$ (right). The gray surfaces are the obstacles’ walls.
3.3 Tortuosity as the inertia onset indicator in simple cubic systems
@@
We continue our analysis by looking at the Reynolds number dependence of $T_{\Omega}$ and $T_{s}$ for the SC sample, shown in Fig. 16. In the Darcy regime ($Re_{k}\rightarrow 0$), the values of both tortuosities stay constant and are compliant with each other, as shown in [27]. We attribute small visible discrepancies between the two tortuosities to the errors introduced during the generation of streamlines and the identification of percolating/recirculation volumes. When the Reynolds number increases, the streamline-based tortuosity, $T_{s}$, decreases for both systems, which was observed in stochastic and periodic porous media by other authors [22, 15]. On the other hand, when the volume-based tortuosity, $T_{\Omega}$, is concerned, the inertial effects manifest themselves differently depending on the porosity, similarly to the stochastic porous media (cf. Fig. 11). In the lower-porosity case, its value decreases slightly in the range $Re_{k}\in[5\cdot 10^{-4};3\cdot 10^{-3}]$ to later increase significantly until the steady-state regime ends. Such behavior resembles the one observed for the lower-porosity stochastic porous media in Fig. 11 and also for a real-life porous sample [16]. For the higher-porosity sample, the volume-based tortuosity follows the streamline-based one closely and diverges from it only at $Re_{k}\approx 10^{3}$. Similarly to the lower-porosity system, the initial decline of $T_{s}$ is followed by its rise, to a much smaller extent, however. This is in line with the behavior of $T_{\Omega}$ for the higher-porosity stochastic porous media shown in Fig. 11, and in [14].
@@
The strictly decreasing behavior of the streamline-based tortuosity can be attributed to the straightening of the percolating streamlines, which follows the growth of the recirculation zones volumes in the cavities and obstacles wakes. The volume-integrated tortuosity exhibits a more complex behavior, dependent on the sample’s porosity. Eq. (14) shows that the difference between $T_{\Omega}$ and $T_{s}$ is related to the integral of velocity magnitude inside the recirculation zones. Thus, we will further investigate the fall and the rise of the volume-based tortuosity in terms of the volume of the recirculation zones and the field of the velocity magnitude therein.
@@
3.4 Origin of the $T_{\Omega}$ decrease and rise – volume of the recirculation zones
Let us first focus on the numerator of the recirculation volume-related term from Eq. (14), $I[V]_{v}/I[V_{0}]_{p}$. For convenience, we divide it by the total fluid volume, $I[1]$, i.e.